OpenStudy (anonymous):

What three consecutive odd numbers have a sum of 45? WILL FAN AND MEDAL!

OpenStudy (blyue):

I got the numbers 13, 15, and 17. :) There's probably a more complicated way to solve this out, but... here's what I did. 45 divided by 3 That's 15. 15 + 15 + 15 = 45 To make them consecutive, subtract 2 from one of the 15s and add it to another. (15 - 2) + 15 + (15 + 2) 13 + 15 + 17 = 45 :)

OpenStudy (anonymous):

Oh wow thx ^-^

OpenStudy (blyue):

You're very welcome. :)

OpenStudy (anonymous):

let the numbers be n+1,n+3,n+5 n+1+n+3+n+5=45 3n+9=45 3n=45-9=36 n=36/3=12 so numbers are 12+1,12+3,12+5 or 13,15,17